if r= 0.8,the variation in dependent variable Y due to independent variable X is about?
Answer by math_tutor2020(3817)


r = 0.8
r^2 = (0.8)^2
r^2 = 0.64

64% of the variation in the dependent variable Y is explained by the variance in the independent variable X.

Answer: 64%